OpenStudy (anonymous):

If 3 cows can eat 16 bales of hay in 8 days, how long will it take 9 cows to eat 7 bales of hay?

OpenStudy (dumbcow):

16/8 = 2/1 3 cows eat at a rate of 2 bales per day multiply by 3 9 cows then eat at rate of 6 bales per day set up proportion using daily rate \[\frac{7}{d} = \frac{6}{1}\] \[d = \frac{7}{6}\] It take 9 cows (7/6) days to eat 7 bales of hay